# More Modality, More AI: Exploring Design Opportunities of AI-Based Multi-modal Remote Monitoring Technologies for Early Detection of Mental Health Sequelae in Youth Concussion Patients

### v1

@@ 3.2.3. Clinicians’ Challenges in Communicating with Patients and Families About Mental Health Issues
Communication challenges persist not only outside the clinics but also during in-person consultations. During in-person visits, parents often accompany their children. However, some patients may feel uncomfortable to disclose sensitive mental health issues. One clinician (P1) described a case where a patient with severe anxiety disclosed self-harm and suicidal thoughts only after requesting to speak privately, away from their parents’ presence.
@@
“I stepped out of the I had the mom stay in the room……And I said, ’Is there something wrong? Is there something going on?’ He said, ’Yeah, I need to talk to you privately.’” (P1)
@@
Clinicians suspect that a major reason that patients choose not to disclose their mental health status is their fear of being labeled as having a mental illness. Being labeled as having a mental illness can affect how they are perceived socially, which could potentially isolate them from their peers. Some patients may also struggle with the societal expectation of their physical recovery, especially if youth patients are athletes who are expected to perform at a high level of competition.
This highlights the sensitivity of mental health disclosures is intertwined with many societal factors that may not be directly related to mental health symptoms.
@@
Moreover, some concussion clinicians (P1, P4, P5) reported that sometimes both youth concussion patients and their parents may deliberately withhold information about the patient’s mental health. In one case (P4), a youth concussion patient was a successful athlete, and the parents downplayed their child’s mental health symptoms during clinical visits. The parents may expect to expedite the child’s recovery process so that they can return to their athletic career sooner. These parents may worry that prolonged rest due to mental health sequelae could hinder their child’s athletic performance and career trajectory. However, when mental health issues are not disclosed and addressed in a timely manner, the issues could get worse and lead to serious consequences, such as self-harm or an increased risk of suicidal behavior. The consequences of severe mental health issues could lead to long-term physical and cognitive impairment of the patients as well as hinder an athlete’s ability to return to sports.
@@
”Their (patients) parents are very vested in the child’s athletic career……the parents come in and there’s obviously an agenda to minimize their symptoms. Or you know, just try not to buck the system, if it’s gonna limit their (patients) activity.” (P4)
@@
To detect undisclosed mental health issues, clinicians rely on a combination of strategies. One approach is to build trust and collaborate directly with parents. Another is to observe the patient’s behavior during consultations. For example, clinicians (P1, P4) can observe patients’ behaviors, such as signs of anxiety or a persistently quiet and disengaged attitude, during clinical visits. Moreover, worsened concussion symptoms such as insomnia and low activity levels may indicate underlying issues beyond the concussion itself, such as mental health sequelae. One clinician (P3) stated that if a youth patient’s concussion symptoms do not improve within four weeks after a concussion, it may indicate underlying mental health issues that are contributing to a prolonged recovery period. However, both observing patients during in-person visits and asking about their sleep and activity levels at home require clinicians to gather this information during the clinical visit. Once the patient returns home, clinicians can hardly access this information.
@@
”Because we expect 96 percent of concussions to be healed by the four-week point. And so if it’s not healed yet, then it’s probably not the concussion, right? It’s probably either a headache syndrome or a mental health concern.” (P3)
3.2.4. Clinicians’ Challenges in Assessing Patients’ Compliance with Clinical Recommendations
At the end of each clinical visit, concussion clinicians typically provide patients with recommendations related to physical activity or sleep. These clinical recommendations are intended to serve as the concussion treatment plan to support both physical and mental recovery. However, understanding whether patients adhere to these recommendations is a persistent challenge.
@@ 3.2.4. Clinicians’ Challenges in Assessing Patients’ Compliance with Clinical Recommendations
Concussion clinicians (P2, P4) rely on patients’ self-reports to assess compliance during the time period between the last visit and the current one, despite these reports are often unreliable. For instance, some patients may claim to have followed the recommendations, but their responses to mental health questionnaires reveal high levels of anxiety or depression, indicating limited or no improvement.
@@
“What we recommended, if they(youth patients) did it or not, they might not have done it… … We don’t know if they’re following them or not. Or they come in, and they say they did, but we don’t know if they did.” (P2)
@@
Concussion clinicians (P) noted that there is currently no effective way to track patients’ compliance with clinical recommendations outside the clinic. Clinicians (P4) mentioned that they often ask patients’ parents about youth patients’ adherence to recommendations but found that information from family members is also hardly reliable. One reason could because some parents are frequently away from home due to work commitments, which leaves them with limited knowledge of patients’ daily health-related information. In such cases, clinicians may reach out to other family members for a closer understanding of the patient’s mental health conditions when they are at home. However, making such phone calls consumes a considerable amount of time in clinicians’ busy schedules, and the information collected through these calls is often limited.
@@
“So I think sometimes the parent comes in with a child who isn’t around them that much at home. Like maybe it’s the working parent, and they’re just sort of the driver and they don’t have a lot of information…… So sometimes I can’t get good information ’cause they’re not engaged as much as I would, you know, would hope.” (P4)
4. Initial System Design
In our formative study, We found that clinicians have a need for remotely monitor the health-related information of youth concussion patients at home, particularly regarding their mental health. Additionally, we found that clinicians face challenges when making mental health-related decisions, particularly during clinical visits, where it can be difficult to gather mental health information from patients and their families through communication. However, clinicians rely on this information to guide their clinical decisions and treatment plans.
Building on these identified challenges and needs, we proposed AI-RPM technology that could potentially assist clinicians in remotely monitoring patients’ at-home data and addressing decision-making difficulties. We then visualized a possible system interface based on these technologies and detailed the specific functions of the system. The goal of our system design is to explore whether concussion clinicians find the technologies useful in remotely collecting youth patient mental health-related information outside of clinics and supporting clinicians’ challenges associated with mental health-related decision-making.
4.1. Applicable AI-RPM Technology
4.1.1. Wearables in Tracking Patients’ Sleep Pattern and Physical Activities
@@
Based on Section 3.2.2, clinicians expressed a desire to monitor patients’ at-home quantitative data, such as sleep and exercise patterns, as these factors are closely related to concussion recovery and mental health status. However, they currently lack effective ways to access real-time health data from patients at home. In addition, as noted in Section 3.2.4, clinicians frequently ask patients whether they adhere to clinical recommendations, such as maintaining proper sleep and exercise routines. However, clinicians often struggle to verify the accuracy of patients’ self-reported information.
@@
To address these challenges, we propose that wearable devices such as smartwatches could help clinicians collect real-time sleep and exercise data from patients and provide these data to clinicians for timely review. Wearables offer a non-invasive and natural way to monitor patients’ health metrics without significantly disrupting their daily routines. As discussed in Section 2.2, wearable devices can continuously monitor patients’ sleep and physical activity data in real-time (Dias and Paulo Silva Cunha, 2018; Baig et al., 2017; Bate et al., 2023). We believe that patients’ sleep patterns and physical activity data collected through wearables could enable clinicians to remotely access at-home patient data, track adherence to sleep and exercise recommendations, and ultimately make more informed mental health-related clinical decisions.
4.1.2. LLM-Powered Conversational Agents in Supporting Patients Self-Report Mental Health Symptoms
@@
In Section 3.2.2, clinicians expressed a need to understand patients’ experiences with symptoms such as panic attacks, including what specifically triggered them. However, patients often struggle to accurately recall these details, and such subjective experiences cannot be easily quantified using standardized scales or medical instruments. Additionally, based on Section 3.2.3, concussion clinicians reported that youth concussion patients sometimes conceal their mental health issues from their parents. It was only during a private conversation with the clinician that the patient disclosed the truth about their mental health issues.
@@
Given this challenge, we hypothesize that LLM-powered conversational agents (CAs) (Yang et al., 2024; Mahmood et al., 2023; Chan et al., 2024) could help clinicians collect mental health-related information effectively. CAs can be deployed in patients’ homes to engage in natural language conversations, ask relevant questions, and record dialogue content. AI can then analyze historical conversations to summarize the content of the conversation, which allows clinicians to quickly gain insights into their mental well-being. One study found that children perceive conversational agents as warmer and more reliable than their parents (Van Brummelen et al., 2023). Therefore, we speculate that CAs could encourage patients to disclose their mental health issues, enable AI to analyze these conversations, extract key information, and securely transmit relevant insights to clinicians while preserving privacy. As a result, clinicians could review patients’ historical conversations with CAs, gaining insights into their subjective experiences, emotional states, and mental health-related life events—ultimately supporting a more accurate assessment of patients’ mental health.
4.1.3. AI Risk Prediction in Supporting Mental Health Sequelae Detection
@@
As discussed in Section 3.2.3, clinicians reported that patients and their parents sometimes deliberately conceal mental health conditions, making it difficult for clinicians to develop appropriate treatment plans. In addition, clinicians noted that the severity of mental health issues in a patient influences treatment plans and subsequent clinical decisions by clinicians, such as whether a referral is necessary. Given these two challenges, we believe that AI-powered risk prediction for disease progression could provide valuable support to clinicians. AI has the capability to analyze patients’ health data and use algorithms to predict the likelihood of developing a condition over a certain period (Collins and Moons, 2019). Recent studies have already explored AI-driven risk prediction for mental health sequelae following a concussion and have found that AI predictions can achieve high accuracy (Dabek et al., 2022). By leveraging AI risk score prediction, clinicians may no longer have to rely solely on patients’ self-reported mental health status and daily behaviors when making assessments. Moreover, if patients or their parents choose to withhold information, AI could help clinicians detect potential mental health sequelae in a timely manner. If a patient is flagged as having a high risk of developing mental health sequelae, the clinician can proactively schedule follow-up visits and make necessary mental health-related clinical decisions. Therefore, we hypothesize that AI-driven risk prediction has the potential to assist concussion clinicians in detecting mental health sequelae more effectively.
@@
Figure 1. The initial system design draft for evaluation and gathering information from concussion clinicians: (A) AI risk score prediction for mental health sequelae, (B) patient self-report symptoms, (C) patient sleep and physical activities.
4.2. Preliminary UI Design of the System
Based on the potentially applicable technologies that could support concussion clinicians in addressing these challenges, we designed three modules for our initial system design, as shown in Figure 1: (1) AI risk score prediction for mental health sequelae (Fig. 1.A), (2) patient self-reported symptoms (Fig. 1.B), and (3) patient sleep and physical activity tracking (Fig. 1.C). During our UI design process, we adhered to Human-Centered AI (HCAI) principles to ensure a user-friendly interface. Our goal with this design is to visualize the system architecture and interface based on the previously proposed technologies, which makes our system design more intuitive for clinicians. Then, we can use the initial system design to gather clinicians’ perceptions and feedback on visual representation, data types, and technologies.
@@ 4.2.1. The AI Risk Score Prediction Module
The first module is AI Risk Score Prediction(Fig 1.A). The module shows the probability of mental health sequelae within one year with a range from 0% to 100%, where the higher probability score signifies a higher probability of experiencing mental health sequelae for clinicians to review. Feature importance scores are also provided to provide additional explanations for clinicians to decide whether to trust or not to trust the AI-predicted score. This module also offers text-based explanations to help clinicians understand the sources of the factors that contribute to the risk score. At the bottom is the risk score chart. Clinicians can hover over any day on the risk score chart to view the likelihood that young patients develop mental health sequelae at that specific point in time. Additionally, clinicians can select different dates to review patient data from various time periods. This module aims to enable clinicians to remotely monitor youth patients’ mental health sequelae without requiring in-person consultations, which supports concussion clinicians in making timely, informed decisions based on patients’ mental health conditions.
4.2.2. The Self-Report Symptoms Module
The second module is the self-reported symptoms (Figure 1.B). On the right side of the module, a history of conversations between youth concussion patients and conversational agents (CAs) is available for concussion clinicians to review. Based on the conversation history, AI detects the symptoms that are mentioned by patients and highlights categories that are relevant to mental health sequelae, such as anxiety, depression, and suicidal ideation. The categories are displayed on the left side of the panel. If a patient self-reports symptoms within a specific category, the corresponding indicator changes from green to red, which reminds clinicians to review the entry. To facilitate efficient navigation, clinicians can click on a category name or use the search bar to enter keywords, which prompts the right-side module to display the relevant conversation excerpts. Additionally, clinicians can select different dates to review patient data from various time periods. The primary goal of this module is to allow concussion clinicians to quickly assess patients’ mental health symptoms at home, extending support beyond clinical settings.
4.2.3. The Sleep and Physical Activities Module
Lastly, we designed the sleep and physical activities module (Fig. 1.B). The module presents sleep-related information that concussion clinicians want to track for their patients at home, including sleep hours, light sleep, and wake-up time, as well as activity-related data such as step count. When a clinician selects any category, the module displays the patient’s relevant data from the past week. Clinicians can also select different dates to review patient data from other time periods. Additionally, the module provides the source of the data, which is through wearable sensors, such as smartwatches. The goal of this module is to assist concussion clinicians in monitoring youth concussion patients’ sleep and physical activity at home, addressing the current gap in existing clinical systems that lack this information.
5. Design Interface Evaluation
In the previous section, we introduced the potential of AI-RPM technologies to assist clinicians in remotely collecting patient information and supporting mental health-related decision-making. We then designed a system that integrates these technologies, focusing on how the collected data is presented and utilized. Based on the initial system design, we conducted a design interface evaluation (Savage, 1996) to gather feedback from the six original clinicians who represent our target end-users. The goal of this section is to explore concussion clinicians’ perceptions of the system—specifically, whether they find these technologies useful for remote patient monitoring and how the system might better support them in addressing challenges related to mental health decision-making.

### v2

@@ 3.2.3. Clinicians’ Struggle to Communicate Mental Health Issues with Patients and Their Families
Communication challenges persist not only outside the clinics but also during in-person consultations. During in-person visits, parents often accompany their children. However, some youth concussion patients feel uncomfortable disclosing sensitive mental health issues with their parents present. P1 described a case where a patient with severe anxiety disclosed self-harm and suicidal thoughts only after requesting to speak privately, away from their parents’ presence.
@@
“I stepped out of the [room], I had the mom stay in the room…… and I said, ’Is there something wrong? Is there something going on?’ He said, ’Yeah, I need to talk to you privately.’ ” (P1)
@@
Clinicians suspect that a major reason for this is patients’ fear of mental health stigma (Association, 2023). The label of having a mental illness can affect how patients are perceived socially, which could potentially isolate them from their peers. Another reason not to disclose mental health issues is the societal expectation toward youth concussion patients. Specifically, if youth patients are athletes who are expected to perform at a high level of competition, they may struggle with the societal expectation of their physical recovery.
@@
Not only do patients withhold their mental health condition, but sometimes their parents also deliberately withhold information about the patient’s mental health condition as reported by concussion clinicians (P1, P4, P5). In one case encountered by P4, the youth concussion patient was a successful athlete, and the parents downplayed their child’s mental health symptoms during clinical visits. These parents may worry that prolonged rest due to mental health sequelae could hinder their child’s athletic performance and career trajectory. So the parents downplayed their child’s mental health symptoms, hoping for their child to return to their athletic career sooner. However, when mental health issues are not disclosed and addressed in a timely manner, the issues could get worse and lead to severe mental health issues, such as self-harm or an increased risk of suicidal behavior. The consequences of severe mental health issues could lead to long-term physical and cognitive impairment of the patients, which might eventually hinder an athlete’s ability to return to sports in their lives.
@@
“Their (patients) parents are very vested in the child’s athletic career……the parents come in and there’s obviously an agenda to minimize their symptoms. Or you know, just try not to buck the system, if it’s gonna limit their (patients) activity.” (P4)
@@
To detect undisclosed mental health issues, clinicians rely on a combination of strategies. One approach is to build trust and collaborate directly with parents. Another is to observe the patient’s behavior during consultations. For example, clinicians (P1, P4) observe patients’ behaviors, such as signs of anxiety or a persistently quiet and disengaged attitude, during clinical visits. Moreover, worsened concussion symptoms such as insomnia and low activity levels may indicate underlying issues beyond the concussion itself, such as mental health sequelae. P3 stated that if a youth patient’s concussion symptoms do not improve within four weeks after a concussion, it may indicate underlying mental health issues that are contributing to a prolonged recovery period. However, observing patients’ behavior or asking about their sleep patterns requires patients to be physically present during the clinical visit. Once the patients return home, clinicians cannot access this information.
@@
“Because we expect 96% of concussions to be healed by the four-week point. And so if it’s not healed yet, then it’s probably not the concussion, right? It’s probably either a headache syndrome or a mental health concern.” (P3)
3.2.4. Clinicians’ Struggle to Evaluate Patients’ Compliance Beyond the Clinic
@@
At the end of each clinical visit, concussion clinicians typically provide patients with recommendations related to physical activity levels or sleep schedules. These clinical recommendations are intended to serve as the concussion treatment plan to support both physical and mental recovery after a concussion. However, understanding whether patients adhere to these recommendations is a persistent challenge. Although concussion clinicians (P2, P4) can rely on patients’ self-reports to assess compliance during the time period between the last visit and the current one, these reports are often unreliable. For instance, some patients may claim to have followed the recommendations, but their responses to mental health questionnaires reveal high levels of anxiety or depression, indicating limited or no improvement.
@@
“What we recommended, if they (youth patients) [said they] did it or not, they might not have done it… … We don’t know if they’re following them or not. Or they come in, and they say they did, but we don’t know if they did.” (P2)
@@
However, there is currently no effective way to track patients’ compliance outside the clinic. P4 mentioned that they often ask patients’ parents about their children’s adherence to the recommendations, but the information provided by the parents may also be unreliable. One reason could be that some parents are frequently away from home due to work commitments, which leaves them with limited knowledge of patients’ daily health-related information. In such cases, clinicians may reach out to other family members by phone for a closer understanding of the patient’s mental health issues when they are at home. However, making such calls consumes a considerable amount of time in clinicians’ busy schedules, and the information collected through these calls is often limited.
@@
“I think sometimes the parent comes in with a child who isn’t around them (parents) that much at home. Like maybe it’s the working parent, and they’re just sort of the driver, and they don’t have a lot of information…… Sometimes I can’t get good information cause they (parents) are not engaged as much as I would, you know, would hope.” (P4)
4. Preliminary System Design
In the formative study, we identified clinicians’ urgent need to remotely monitor the mental health of youth concussion patients at home, as well as the challenges they face in clinical decision-making. In this section, we explored the design opportunities of AI-RPM technologies in supporting clinicians’ needs and decision-making during their practice. Subsequently, we proposed a preliminary system design (Fig. 2) based on the AI-RPM technologies with detailed features. The goal of the design is to explore whether concussion clinicians find the AI-RPM system useful for supporting remote monitoring and decision-making, how the system can be improved, and how they would use it in practice.
4.1. Design Considerations for AI-RPM Systems
4.1.1. Wearables for Tracking Patient Physiological Data
@@
Based on Section 3.2.2, clinicians expressed a desire to monitor patients’ at-home quantitative data, such as sleep and exercise patterns, as these factors are closely related to concussion recovery and mental health status. In addition, as noted in Section 3.2.4, clinicians frequently ask patients about their clinical recommendation compliance, such as maintaining proper sleep and exercise routines. However, clinicians currently lack effective ways to access these at-home data and often struggle to verify the accuracy of patients’ self-reported information during clinical visits.
@@
To address these challenges, wearable devices such as smartwatches can be used to help clinicians collect real-time sleep and exercise data from patients (Dias and Paulo Silva Cunha, 2018; Baig et al., 2017; Bate et al., 2023). Wearables offer a non-invasive and natural way to monitor patients’ health metrics without significantly disrupting their daily routines. Moreover, AI has the potential to process the massive amount of data collected from wearables and provide these data to clinicians for timely review. We believe that with access to at-home patient data, clinicians can become aware of patients’ health conditions in real-time, and track patients’ compliance with sleep and exercise recommendations, thus ultimately having the ability to make more informed mental health-related clinical decisions, such as setting up a follow-up meeting for further assessment.
4.1.2. LLM-based CAs for Patient Self-Reporting in Mental Health
@@
In Section 3.2.2, clinicians emphasized the need for a comprehensive understanding of patients’ experiences with mental health symptoms, such as panic attacks, to provide effective care. However, patients often struggle to accurately recall these details, and such subjective experiences cannot be easily quantified using standardized scales or medical instruments. Additionally, based on Section 3.2.3, concussion clinicians reported that youth concussion patients sometimes conceal their mental health issues from their parents. It was only during a private conversation with the clinician that the patient disclosed their mental health status.
@@
Given this challenge, we hypothesize that LLM-based CAs (Yang et al., 2024; Mahmood et al., 2023; Chan et al., 2024) could help clinicians collect mental health-related information effectively. CAs can be deployed in patients’ homes to engage in natural language conversations, ask relevant questions, and record dialogue content. AI can then analyze historical conversations to summarize the key information of the conversation, such as their patient’s suicidal intentions. From the perspective of youth patients, the deployment of a CA in their rooms could increase the chance of disclosing their mental health issues, if there are any. One study found that children perceive conversational agents as warmer and more reliable than their parents (Van Brummelen et al., 2023). From the perspective of clinicians, gaining such key information allows them to review patients’ historical conversations with CAs, gain accurate information about patients’ mental health-related experiences while preserving privacy, and conduct interventions if necessary. Therefore, we speculate that CAs have the potential to act as a communication bridge between clinicians in the hospital and patients at home and provide patients with self-reported symptoms. Ultimately, CAs could support clinicians in conducting a more accurate assessment of patients’ mental health and making timely decisions.
4.1.3. AI-based Risk Prediction for Detecting Mental Health Sequelae
@@
As we have indicated in Section 4.1.2, clinicians reported that patients and their parents sometimes deliberately conceal mental health conditions. In addition, clinicians noted that the severity of mental health symptoms in a patient influences their subsequent clinical decisions, such as treatment plans or providing a referral. However, evaluating the severity of mental health symptoms remains challenging. Given these two challenges, we believe that AI risk prediction for disease progression could provide valuable support to clinicians in our setting. AI has the capability to analyze patients’ health data from the EHR system and utilize algorithms to predict the likelihood of developing specific conditions over a given period (Collins and Moons, 2019). In particular, recent studies have already explored AI-driven risk prediction for mental health sequelae following a concussion and have found that AI predictions can attain high accuracy (Dabek et al., 2022). By leveraging AI risk score prediction, clinicians may no longer have to rely solely on the conversations with patients during clinical visits to make mental health assessments. If patients or their parents choose to withhold information, AI could help clinicians detect potential mental health sequelae in a timely manner. Moreover, the severity of the mental health symptoms is normally associated with the AI prediction score. The AI prediction score can help clinicians better assess the severity of a patient’s mental health symptoms and make informed decisions accordingly.
@@
Figure 2. The preliminary system design based on concussion clinicians’ challenges and needs: (A) AI risk score prediction for mental health sequelae, (B) patient self-report symptoms, (C) patient sleep and physical activities.
4.2. Three Modules of Preliminary System Design Based on the Design Considerations
Based on the potentially applicable technologies that could support concussion clinicians in addressing their challenges and supporting their decision-making, we designed three modules for our preliminary system, as shown in Fig. 2. We described each module in detail in this section.
4.2.1. The Sleep and Physical Activities Module
The sleep and physical activities module (Fig. 2 B) presents sleep (sleep hours, light sleep, and wake-up time) and activity data (step count) that clinicians want to track. These data can be collected by wearables as we discussed on Section 4.1.1 Moreover, clinicians can select any category to review the patient’s relevant data from the past week. Clinicians can also select different dates to review patient data from other time periods. Additionally, the module provides the source of the data for increasing clinicians’ understanding of the module and their trust towards the system.
4.2.2. The Self-Report Symptoms Module
To display the patients’ self-reported symptoms at home, which were collected by CAs as we proposed in Section 4.1.2, we designed the self-reported symptoms (Fig. 2 B). On the right side, a history of conversations between youth concussion patients and CAs is available for concussion clinicians to review. Based on the conversation history, AI detects the symptoms that are mentioned by patients and highlights key information that are relevant to mental health sequelae, such as anxiety, depression, and suicidal ideation. If a patient self-reports symptoms within a specific category, the corresponding indicator changes from green to red. We use the color red to catch clinicians’ attention. To facilitate efficient navigation, clinicians can click on a category name or use the search bar to enter keywords, which prompts the right-side module to display the relevant conversation excerpts. Similar as the sleep and physical activities module, time selection lets clinicians view data from different periods.
@@ 4.2.3. The AI Risk Score Prediction Module
The AI Risk Score Prediction (Fig. 2 A) module leverages AI’s ability that we discussed in Section 4.1.3 to show the probability of mental health sequelae. The prediction timeframe is within one year with a range from 0% to 100%, where the higher probability score signifies a higher probability of experiencing mental health sequelae for clinicians to review. Moreover, feature importance is used to provide data that influence the AI prediction score, which allows clinicians to decide whether to trust the AI-predicted score or not. To increase AI explanations and transparency, we provide text-based explanations to help clinicians understand the risk score and its source. At the bottom is the risk score chart. Clinicians can hover over any day on the chart to view the likelihood of developing mental health sequelae at that specific point in time. Additionally, clinicians can select different dates to review patient data from various time periods.
5. Evaluation Study of the Preliminary System Design
To evaluate the usefulness of our preliminary design, we conducted an evaluation study. The goal of the evaluation study is to gather clinicians’ feedback and insights on our preliminary designs with different AI-RPM technologies. Specifically, we want to explore (1) whether they find these AI-RPM technologies useful for remotely monitoring mental health-related information of youth concussion patients and (2) how to revise the preliminary system design from data type, visualization, interaction, and usability perspectives to better support clinicians in addressing challenges and decision-making in their workflow.